Two ambitious business graduates, Boris and Isabelle, are considering purchasing Premier Pizza, a frozen pizza manufacturer. Forecasted annual sales for the coming year are $10.4 million, with operating costs equal to 79% of sales, depreciation is 3% of sales and required annual investment in equipment is 3% of sales. Sales, costs, and investments are expected to grow 5% in perpetuity. On the basis of their analysis of the industry, Boris and Isabelle anticipate financing their company with 21% debt. The required rate of return on the debt will be 5% and the required rate of return on the equity will be 18%. Premier Pizza’s corporate tax rate is 35%. The risk-free interest rate is 3% and the market risk premium is 8%.


a. What is the maximum price Boris and Isabelle should be willing to pay for Premier Pizza? (Round your answer to 2 decimal places.)
Maximum price   $   million


b. A national grocery chain, Fresh Foods, is also considering making an offer for Premier Pizza. The levered equity beta of the grocery chain is 0.9, its cost of debt is 4%, it is 44% debt-financed, and it has a 35% tax rate. What is the maximum price that Fresh Foods should be willing to pay for Premier Pizza? Assume that Fresh Foods’s forecast for Premier’s cash flows is the same as Boris and Isabelle’s. (Round your answer to 2 decimal places.)
Maximum price    $   million
Answer #1

a) Based on information given, the Free cashflow and Cost of capital is worked out below. Thereafter the value of firm is worked out.

Free Cashflow Computation Amt in million $
Sales (a)                                     10.400
Operating Cost @ 79% of Sales (b)                                       8.216
Operating Profit c = (a-b)                                       2.184
Less: Tax @ 35% of Operating Profit ( d)                                       0.764
Add: Tax benefit on Depreciation = (35% on Depreciation) ( e )                                       0.109
Operating Cash Flow f =(c-d+f)                                       1.529
Investment in equipments @ 3% of Sales ( g )                                       0.312
Free Cashflow (FCF) h = (f-g)                                  1.21680
Note: Depreciation @ 3% on Sales (3% on $ 10.40 million)                                         0.312
Tax benefit @ 35% on Depreciation (35% on $ 0.312 million) 0.1092
Cost of Capital Computation
Cost of Debt (kd)= Interest rate*(1-tax rate) = 5%*(1-35%) 3.25%
Cost of Equity (ke) 18.00%
Weights
Weight of Debt (wd) 21%
Weight of Equity (we) 79%
WACC = (kd x wd)+(ke x wd) 14.9025%
It is given that Sales, Cost and Investment will grow at 5%, hence g= 5%
Value of firm = FCF/(WACC-g)
= 1.2168/(14.9025%-5%)= 1.2168/(9.9025%)
= $ 12.29 million

The maximum price,x that Boris and Isabelle will give for premier pizza is $12.29 million.

b) Value of firm that Fresh foods will pay

Cost of Capital for Fresh Foods, a national grocery chain is as follows

Cost of Capital Computation for Fresh foods
Cost of Debt (kd)= Interest rate*(1-tax rate) = 4%*(1-35%) 2.60%
Cost of Equity (ke) = Riskfree rate+Beta of grocery chain*(Market Risk premium) = 3%+0.9*8% 10.20%
Weights
Weight of Debt (wd) 44%
Weight of Equity (we) 56%
WACC = (kd x wd)+(ke x wd) 6.8560%
As per info FCF is same for Fresh food and Boris & Isabelle's, hence value of firm would be
It is given that Sales, Cost and Investment will grow at 5%, hence g= 5%
Value of firm = FCF/(WACC-g)
1.2168/(6.860%-5%) = 1.2168/(1.860%)
= $ 65.42 million

Maximum price that Fresh foods will give for Premier pizza is = $65.42 million